Solve the following application problem (show equation and then solution) Round answers to the nearest thousandths and properly label answer.
The length of a rectangle is 5cm more than it's width. If the area of the rectangle is 90cm^2 find the dimensions.


Answer by mananth(16946)

let the width be x cm
the length will be x+5 cm
..
Area of rectangle = L*W
..
x(x+5)=90
x^2+5x=90
x^2+5x-90=0
x^2+15x-6x-90=0
x(x+15)-6(x+15)=0
(x+15)(x-6)=0
x+-15 or 6.
The width cannot be negative
so width = 6 cm.
Length will be 6+5=11cm